A spaceship is traveling at half the speed of light (0.5c). Their onboard clock shows that they traveled for a full year. How much time passed on Earth during their “year”? (Give your answer in years)


1
Expert's answer
2022-05-06T15:34:17-0400

"t=\frac{t_0}{\sqrt{1-\frac{v^2}{c^2}}}=1.15^y."